# Heterogeneous Intertemporal Treatment Effects via Dynamic Panel Data Models

### v1

@@ 5 Conclusion
$=$ $E\left[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))\left|\begin{array}[]{c}\phi_{2}(y_{0},y_{1},d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},\\ \phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},\text{ \ }Y_{0}=y_{0}\end{array}\right.\right]$
$=$ $E[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))=y_{1},Y_{0}=y_{0}]\text{,}$
@@
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}(d_{2}))|Y_{0}=y_{0}]\text{,}$
where the last equality holds because of (33) and the fact that $Y_{1}(d_{1})$ is a function of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$. Again, the right-hand side does not vary with $d_{1}$.
@@
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|\phi_{1}(y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))|Y_{0}=y_{0}]\text{,}$
@@
$=$ $E\left[\lambda_{2}(0,0,y_{0},Y_{1}(0),\alpha(0),\varepsilon_{2}(0))-Y_{1}(0)\left|\begin{array}[]{c}\phi_{2}(y_{0},Y_{1}(d_{1}),d_{1},\xi^{1}(d_{1}),\eta_{2}(\cdot))=d_{2},\\ \phi_{1}(Y_{0},\xi_{0},\eta_{1}(\cdot))=d_{1},Y_{0}=y_{0}\end{array}\right.\right]$
where $Y_{2}(0,0)-Y_{1}(0)$ is only a deterministic function of $(Y_{0},\alpha(0),\varepsilon_{1}(0),\varepsilon_{2}(0))$. Under (33), the right-hand side does generally vary with $(d_{1},d_{2})$ even after controlling for $Y_{0}$. This is because the event “$(D_{1},D_{2})=(d_{1},d_{2})$” does restrict the joint support of $(Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}),\xi^{1}(d_{1}),\eta^{2}(\cdot))$. [^8] Note $Y_{1}(d_{1})=\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$ affects the definition of conditioning events due to its entry through $\phi_{2}(\cdot)$. Hence, even after conditioning on $Y_{0}=y_{0}$, the right-hand side is in general non-degenerate in $(d_{1},d_{2})$.
Appendix B: Consistency of Adjusted IPW Estimator
In this appendix, we establish the consistency of the adjusted IPW estimator $\widehat{\mu}_{\tau_{2}}$, which is defined in (20). Maintain the following conditions:
@@
[…]
Condition (57) requires uniform convergence of preliminary estimators $\widehat{M}_{1i},\widehat{M}_{2i}$; it holds for many nonparametric (e.g. Nadaraya-Watson) estimators of conditional means under standard conditions. Condition (58) include standard, mild requirement that the sample moments of $M_{1i},M_{2i}$ are well-behaved.
Corollary B1
[…]
Proof of Corollary B1. By the Mean Value Theorem,
$\hat{\mu}_{\tau_{2}}=\frac{1}{n}\sum\nolimits_{i}\left[\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}+\left(\frac{-1}{\tilde{M}_{1i}}\right)\left(\hat{M}_{2i}-M_{2i}\right)+\left(\frac{\tilde{M}_{2i}}{\tilde{M}_{1i}^{2}}\right)(\hat{M}_{1i}-M_{1i})\right]\text{,}$
where $(\Delta Y_{i2}-M_{2i})/M_{1i}$ is an “infeasible” summand that involves the actual population moments $M_{1i},M_{2i}$, and $\tilde{M}_{ki}$ are intermediate values bounded between $\hat{M}_{ki}$ and $M_{ki}$ for $k=1,2$. By the triangular inequality,
$\left|\hat{\mu}_{\tau_{2}}-\frac{1}{n}\left(\sum\nolimits_{i}\frac{\Delta Y_{i2}-M_{2i}}{M_{1i}}\right)\right|$
$\leq$ $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\left|\hat{M}_{2i}-M_{2i}\right|+\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\left|\hat{M}_{2i}-M_{2i}\right|$
$\leq$ $\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{2i^{\prime}}-M_{2i^{\prime}}\right|\right)+\left(\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\right)\left(\sup_{1\leq i^{\prime}\leq n}\left|\hat{M}_{1i}-M_{1i}\right|\right)\text{.}$
By (57), the two suprema of differences are $o_{p}(1)$, i.e., converge in probability to 00.
Next, we establish that $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$, i.e. is bounded in probability. For this, we want to show “$\forall\varepsilon>0$, $\exists\mathcal{C}_{\varepsilon}>0$ such that $\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}<\varepsilon$ for $n$ large enough.” By construction,
$\Pr\{\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\}\leq\Pr\left\{\max\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|^{-1}>\mathcal{C}_{\varepsilon}\right\}$ (59)
$=$ $\Pr\left\{\min\nolimits_{i\leq n}\left|\tilde{M}_{1i}\right|<1/\mathcal{C}_{\varepsilon}\right\}\leq\Pr\left\{\min\nolimits_{i\leq n}M_{1i}-\tilde{e}_{1n}<1/\mathcal{C}_{\varepsilon}\right\}$
$=$ $\Pr\left\{\tilde{e}_{1n}>\min\nolimits_{i\leq n}M_{1i}-1/\mathcal{C}_{\varepsilon}\right\}\text{,}$
where $\tilde{e}_{kn}\equiv\sup_{i\leq n}\left|\tilde{M}_{ki}-M_{ki}\right|$ for $k=1,2$, and the second inequality is due to triangular inequality and positivity of $M_{i}$. By condition (i) in (58), $\min_{i}M_{i}-1/\mathcal{C}_{\varepsilon}$ is strictly positive when $\mathcal{C}_{\varepsilon}$ is large enough. The uniform convergence of $\hat{M}_{1i}$ in (57) then implies the right-hand side of (59) converges to zero as $n\rightarrow\infty$. This proves $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{1i}\right|^{-1}=O_{p}(1)$.
Finally, we establish $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$ is bounded in probability. Note:
$\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}\leq\left(\frac{1}{n}\sum\nolimits_{i}\left|M_{2i}\right|+\tilde{e}_{2n}\right)\left(\max_{i^{\prime}}\left|\tilde{M}_{1i^{\prime}}\right|^{-2}\right)\text{,}$
due the triangular inequality and definition of $\tilde{e}_{2n}$. The second term on the right-hand side is $O_{p}(1)$ by the same argument above (with $\mathcal{C}_{\varepsilon}$ replaced by $\mathcal{C}_{\varepsilon}^{2}$). By condition (ii) in (58), the stochastic order of the first term on the right-hand side is $O_{p}(1)+o_{p}(1)=O_{p}(1)$. Thus the product of the two terms, and consequently $\frac{1}{n}\sum\nolimits_{i}\left|\tilde{M}_{2i}\right|\left|\tilde{M}_{1i}\right|^{-2}$, is bounded in probability.

### v2

@@ Appendix A Proofs
[…]
$=$ $E[\lambda_{2}(d_{1},d_{2},y_{0},y_{1},\alpha(d_{2}),\varepsilon_{2})|\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})=y_{1},Y_{0}=y_{0}]\text{,}$
@@
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2})|\phi_{1}(y_{0},\xi_{0},\eta_{1})=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{2}(0,0,y_{0},Y_{1}(d_{1}),\alpha(d_{2}),\varepsilon_{2}|Y_{0}=y_{0}]\text{,}$
where the last equality holds because of (35) and the fact that $Y_{1}(d_{1})$ is a function of $(Y_{0},\alpha(d_{1}),\varepsilon_{1})$. Again, the right-hand side does not vary with $d_{1}$.
@@
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})|\phi_{1}(y_{0},\xi_{0},\eta_{1})=d_{1},Y_{0}=y_{0}]$
$=$ $E[\lambda_{1}(d_{1},y_{0},\alpha(d_{1}),\varepsilon_{1})|Y_{0}=y_{0}]\text{,}$
@@
where $Y_{2}(0,0)-Y_{1}(0)$ is a deterministic function of $(Y_{0},\alpha(0),\varepsilon_{1},\varepsilon_{2})$. Under (35), the right-hand side does generally vary with $(d_{1},d_{2})$ even conditional on $Y_{0}$. This is because the event $(D_{1},D_{2})=(d_{1},d_{2})$” does restrict the joint support of $(Y_{0},\alpha(d_{1}),\varepsilon_{1},\xi_{1}(d_{1}),\eta^{2})$. [^12] Note $Y_{1}(d_{1})=\lambda_{1}(d_{1},Y_{0},\alpha(d_{1}),\varepsilon_{1}(d_{1}))$ affects the definition of conditioning events due to its entry through $\phi_{2}(\cdot)$. Hence, even after conditioning on $Y_{0}=y_{0}$, the right-hand side is in general non-degenerate in $(d_{1},d_{2})$.
Appendix B Generalizations
B.1 Identification of Treatment Effects for ARUPO
In this appendix we extend the definition and identification of treatment effects to general sequences of past treatments $d^{t-1}$. That is, for any $d^{t-1}\in\{0,1\}^{t-1}$, generalize (25) to:
$ATT_{t}(d^{t-1},y_{0})\equiv E[\tau_{it}^{*}(d^{t-1})|D_{it}=1,\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
where $\mathcal{H}_{t-1}(d^{t-1},y_{0})\equiv\{D_{i}^{t-1}=d^{t-1},Y_{i0}=y_{0}\}$. We maintain the ARUPO model (Assumption 6), but strengthen Assumption 7 on treatment effects as follows.
Assumption 8
For any initial condition $y_{0}$ and $t\geq 1$,
$E[\tau_{it}^{*}(d^{t-1})|D_{it},\mathcal{H}_{t-1}(d^{t-1},y_{0}),\tau_{i}^{*t-1}(d^{t-2}),\alpha_{i}^{*},\varepsilon_{i}^{*t-1}]=E[\tau_{it}(d^{t-1})|D_{it},\mathcal{H}_{t-1}(d^{t-1},y_{0})],$
where $d^{t-1}\equiv(d^{t-2},d_{t-1})$.
Assumption 8 resembles Assumption 7 except that it also posits the (conditional) serial uncorrelation of past treatment effects. Under Assumption 8,
$E[\tau_{it}(d^{t-1})|D_{it},D_{i}^{t-1}=d^{t-1},Y_{i}^{t-1}]=E[\tau_{it}(d^{t-1})|D_{it},D_{i}^{t-1}=d^{t-1},Y_{i0}],$ (56)
because $Y_{i}^{t-1}$ is a function of $\{Y_{i0},\tau_{i}^{*t-1}(d^{t-2}),\varepsilon_{i}^{*t-1},\alpha_{i}^{*}\}$ conditional on $D_{i}^{t-1}=d^{t-1}$. Extending our IV estimand to the generalized history, we (re)define:
$Z_{it}(d^{t-1},y_{0})\equiv E[X_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
@@ B.1 Identification of Treatment Effects for ARUPO
$\begin{pmatrix}\widetilde{\beta}(d^{t-1},y_{0})\\ \widetilde{\gamma}(d^{t-1},y_{0})\\ \widetilde{\Delta\theta_{t}}(d^{t-1},y_{0})\end{pmatrix}\equiv E[Z_{it}^{\prime}(d^{t-1},y_{0})X_{it}|\mathcal{H}_{t-1}(d^{t-1},y_{0})]^{-1}E[Z_{it}^{\prime}(d^{t-1},y_{0})\Delta Y_{it}|\mathcal{H}_{t-1}(d^{t-1},y_{0})].$ (57)
The next result is analogous to Proposition 6 under Assumption 6 and 7, only with the conditioning event and the causal parameter defined as $\mathcal{H}_{t-1}(d^{t-1},y_{0})$ and $ATT_{t}(d^{t-1},y_{0})$.
Proposition 9
Suppose Assumptions 6 and 8 hold, and:
$E[Z_{it}^{\prime}(d^{t-1},y_{0})Z_{it}(d^{t-1},y_{0})|\mathcal{H}_{t-1}(d^{t-1},y_{0})]$
has full rank. Then the IV slope estimands (57) for $t\geq 3$ recover the causal terms:
$\widetilde{\beta}(d^{t-1},y_{0})$ $=ATT_{t}(d^{t-1},y_{0});$
$\widetilde{\gamma}(d^{t-1},y_{0})$ $=\gamma^{*},$
and the intercept estimand recovers:
$\widetilde{\Delta\theta_{t}}(d^{t-1},y_{0})=\begin{cases}\Delta\theta_{t}^{*}&\text{when $d_{t-1}=0$};\\ \Delta\theta_{t}^{*}+ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0})&\text{when $d_{t-1}=1$.}\end{cases}$
Proof of Proposition 9. Since this proof broadly resembles the proof of Proposition 9, we only note the points of departure and extension.
First, consider the case with $d_{t-1}=0$ so that $\Delta D_{it}=D_{it}$ conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$. The proof of these analogous identification results in Proposition 6 remains mostly the same, with the following changes. First, it conditions on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$ instead of the previous $\mathcal{H}_{t-1}(y_{0})$. Second, $W_{it-1}$ is now a function of $\{\varepsilon_{i}^{*t-1}(\cdot),\alpha_{i}^{*},\tau_{t-2}^{*}(d^{t-2})\}$ once conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$, which requires Assumption 8 to conclude that:
$E[\tau_{it}^{*}(D_{i}^{t-1})|D_{it}=1,W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=ATT_{t}(d^{t-1},y_{0}).$
In the other case, $d_{t-1}=1$ so that $\Delta D_{it}=D_{it}-1$ conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$. Relative to Proposition 6, the conditional mean of the differenced outcome (analogous to (53)) now includes an additional term:
$E[\tau_{it-1}^{*}(D_{i}^{t-2})|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=-ATT_{t-1}(d^{t-2},y_{0}),$
where the equality holds under Assumption 8 because $W_{it-1}$ is a function of $\varepsilon_{i}^{*t-1}(\cdot)$, $\alpha_{i}^{*}$ and $\tau_{t-2}^{*}(d^{t-2})$ once conditional on $\mathcal{H}_{t-1}(d^{t-1},y_{0})$. Plugging in and rearranging yields the linear conditional expectation expression:
$E[\Delta Y_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]=$ $ATT_{t}(d^{t-1},y_{0})E[\Delta D_{it}|W_{it-1},\mathcal{H}_{t-1}(d^{t-1},y_{0})]+$
$\gamma^{*}E[\Delta Y_{it-1}|W_{it-1},\mathcal{H}_{t-1}(y_{0})]+\Delta\theta_{t}^{*}+$
$ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0}).$
The argument for interpreting the OLS estimand as an IV estimand follows as before. Note that in this case, the estimand for the constant term includes an additional intertemporal difference in treatment effects:
$\widetilde{\Delta\theta_{t}}=\Delta\theta_{t}^{*}+ATT_{t}(d^{t-1},y_{0})-ATT_{t-1}(d^{t-2},y_{0}).$